# Uncertainty Quantification in Probabilistic Machine Learning Models: Theory, Methods, and Insights

### v1

@@ Abstract
Index Terms:
Uncertainty quantification, aleatoric, epistemic, Gaussian process latent variable models, predictive distributions.
I Introduction
Uncertainty Quantification (UQ) plays a fundamental role in probabilistic machine learning (ML). ML models are increasingly used in applications that require making decisions and where reliable predictions are essential. This is especially important in fields that demand reliable choices under uncertainty, such as medical diagnosis [1] autonomous systems [2], and scientific modeling in general [3]. However, most traditional ML models provide only point predictions without assessing their confidence. This, in turn, makes it difficult to quantify the reliability of the results [5].
To address this limitation, probabilistic ML models provide a foundational framework for representing uncertainty [4]. By explicitly modeling uncertainty [10], these models not only improve reliability in high-risk applications, but also promote better generalization by distinguishing between known and unknown regions in the input space. Moreover, uncertainty estimates increase model interpretability by clarifying the sources of uncertainty and revealing potential limitations and biases, which emphasizes the role of UQ in characterizing and managing the uncertainty inherent in models and their predictions while providing valuable insights into reliability and confidence calibration. Given the broad impact of UQ, improving uncertainty estimation remains a pressing challenge.
@@ I Introduction
In this paper, we focus on developing a principled framework for estimating both epistemic and aleatoric uncertainty in probabilistic ML models. Since probabilistic ML models comprise a broad range of approaches, we focus on making the problem more tractable by studying the UQ of predictions made by Gaussian process latent variable models (GPLVMs) [9]. Specifically, we explore how uncertainty can be efficiently quantified using scalable approximations, such as Random Fourier Features (RFF)-based Gaussian processes (GPs), which we use to construct GPLVMs.
@@
The paper is organized as follows. First, we provide background on probabilistic ML models in Section II, followed by a brief overview of GPLVMs in Section III. In Section IV, we present the theoretical foundation for evaluating epistemic and aleatoric uncertainties, while in Section V we provide details of the methods used for these evaluations. Section VI discusses experimental results and insights gained from various experiments. Finally, Section VII concludes with final remarks.
II Background
@@
III GPLVMs and inference with GPLVMs
We now briefly introduce GPLVMs with a focus on the uncertainties in their predictions. GPLVMs are powerful tools for modeling high-dimensional data by learning low-dimensional latent representations [9]. They establish a connection between a low-dimensional latent space and a high-dimensional observed space through GPs. Specifically, GPLVMs model data $\bm{Y}\in\mathbb{R}^{N\times d_{y}}$ in a high-dimensional space $\mathcal{Y}$, where $\mathcal{Y}=\mathbb{R}^{d_{y}}$, by assuming that these data are generated from latent variables $\bm{X}$ in a lower-dimensional space, $\mathcal{X}$, with $\bm{X}\in\mathbb{R}^{N\times d_{x}}$ and $\mathcal{X}=\mathbb{R}^{d_{x}}$, where $d_{x}\ll d_{y}$. Here, $N$ represents the number of paired observations between $\mathcal{X}$ and $\mathcal{Y}$.
More specifically, each observed vector $\bm{y}_{n}\in\mathbb{R}^{d_{y}}$, $n=1,2,\ldots,N$ is generated from a corresponding latent vector $\bm{x}_{n}\in\mathbb{R}^{d_{x}}$, where the mapping $\bm{x}_{n}\mapsto\bm{y}_{n}$ is modeled using $d_{y}$ GPs [7]. In other words, for a given latent variable $\bm{x}_{n}$, the observed variable $\bm{y}_{n}$ is drawn according to
@@ III GPLVMs and inference with GPLVMs
Figure 1 presents a visual representation of a GPLVM.
$\bm{x}_{n}$$y_{1,n}$$y_{2,n}$$y_{d_{y},n}$$f_{1}(\bm{x}_{n})$$f_{2}(\bm{x}_{n})$$f_{d_{y}}(\bm{x}_{n})$
Figure 1: A GPLVM with its basic elements, the latent vector $\bm{x}_{n}$ and $d_{y}$ GPs, one for each observed output.
@@
In order to expand our work on UQ to as wide a set of problems as possible, we explore the use of Random Fourier Features (RFF)–based GPs [6]. These GPs approximate an adopted kernel function by a finite-dimensional RFF mapping by
@@
where $\bm{\omega}_{j}$ are random frequencies, which according to Bochner’s theorem [8], are drawn from the power spectral density of the chosen kernel. The functions of interest shown in Fig. 1 are then modeled by
@@
IV Uncertainty quantification: The theory
@@
V Uncertainty quantification: The methods
@@
VI Uncertainty quantification: The insights from experiments
@@
During training, we drew $M=100$ samples from $p(\bm{X}|\bm{Y})$ using the variational distribution of $\bm{X}$. For each test sample, $\bm{y}_{o,*}$ and $\bm{X}^{(m)}$, we generated $L=100$ samples of $\bm{x}_{*}$ according to $\bm{x}_{*}^{(m,\ell)}\sim p(\bm{x}_{*}|\bm{y}_{o,*},\bm{X}^{(m)},\bm{Y})$. We used the generated training samples $\bm{X}^{(m)}$ to compute the epistemic and aleatoric uncertainties defined by (V) and (32), respectively. We conducted experiments where the number of drawn frequencies took values from the set $J\in\{10,25,50,100,200,500\}$. All the frequencies were obtained by sampling from the radial-basis-function kernel.
During evaluation, we computed the aleatoric and epistemic uncertainties as described in the previous section. In each trial, we assumed that one of the output variables, $y_{d,*}$, was missing and had to be predicted using the remaining variables. For example, to predict $y_{1,*}$, we used the observed values $y_{2,*}$, $y_{3,*}$, and $y_{4,*}$ to estimate the latent variable $\bm{x}_{*}$, which was then used to infer $y_{1,*}$. The boxplots in Figs. 2 and 3 summarize average aleatoric and epistemic variances computed on 200 test samples per trial, following training. Each boxplot is constructed from 50 such averages obtained across independent simulations. The circles denote outlier trials. The plots illustrate the consistency of uncertainty estimates and the variability across test instances.

### v2

@@ Abstract
1 Introduction
Uncertainty Quantification (UQ) plays a fundamental role in probabilistic machine learning (ML). ML models are increasingly used in applications that require making decisions and where reliable predictions are essential. This is especially important in fields that demand reliable choices under uncertainty, such as medical diagnosis Abdar et al., (2022) autonomous systems Cheng et al., (2023), and scientific modeling in general Ghanem et al., (2017). However, most traditional ML models provide only point predictions without assessing their confidence. This, in turn, makes it difficult to quantify the reliability of the results Papamarkou et al., (2024).
To address this limitation, probabilistic ML models provide a foundational framework for representing uncertainty Murphy, (2012). By explicitly modeling uncertainty Wang et al., (2025), these models not only improve reliability in high-risk applications, but also promote better generalization by distinguishing between known and unknown regions in the input space. Moreover, uncertainty estimates increase model interpretability by clarifying the sources of uncertainty and revealing potential limitations and biases, which emphasizes the role of UQ in characterizing and managing the uncertainty inherent in models and their predictions while providing valuable insights into reliability and confidence calibration. Given the broad impact of UQ, improving uncertainty estimation remains a pressing challenge.
@@ 1 Introduction
In this paper, we focus on developing a principled framework for estimating both epistemic and aleatoric uncertainty in probabilistic ML models. Since probabilistic ML models comprise a broad range of approaches, we focus on making the problem more tractable by studying the UQ of predictions made by Gaussian process latent variable models (GPLVMs) Titsias and Lawrence, (2010). Specifically, we explore how uncertainty can be efficiently quantified using scalable approximations, such as Random Fourier Features (RFF)-based Gaussian processes (GPs), which we use to construct GPLVMs.
@@
The paper is organized as follows. First, we provide background on probabilistic ML models in Section 2, followed by a brief overview of GPLVMs in Section 3. In Section 4, we present the theoretical foundation for evaluating epistemic and aleatoric uncertainties, while in Section 5 we provide details of the methods used for these evaluations. Section 6 discusses experimental results and insights gained from various experiments. Finally, Section 7 concludes with final remarks.
2 Background
@@
3 GPLVMs and inference with GPLVMs
We now briefly introduce GPLVMs with a focus on the uncertainties in their predictions. GPLVMs are powerful tools for modeling high-dimensional data by learning low-dimensional latent representations Titsias and Lawrence, (2010). They establish a connection between a low-dimensional latent space and a high-dimensional observed space through GPs. Specifically, GPLVMs model data $\bm{Y}\in\mathbb{R}^{N\times d_{y}}$ in a high-dimensional space $\mathcal{Y}$, where $\mathcal{Y}=\mathbb{R}^{d_{y}}$, by assuming that these data are generated from latent variables $\bm{X}$ in a lower-dimensional space, $\mathcal{X}$, with $\bm{X}\in\mathbb{R}^{N\times d_{x}}$ and $\mathcal{X}=\mathbb{R}^{d_{x}}$, where $d_{x}\ll d_{y}$. Here, $N$ represents the number of paired observations between $\mathcal{X}$ and $\mathcal{Y}$.
More specifically, each observed vector $\bm{y}_{n}\in\mathbb{R}^{d_{y}}$, $n=1,2,\ldots,N$ is generated from a corresponding latent vector $\bm{x}_{n}\in\mathbb{R}^{d_{x}}$, where the mapping $\bm{x}_{n}\mapsto\bm{y}_{n}$ is modeled using $d_{y}$ GPs Rasmussen and Williams, (2006). In other words, for a given latent variable $\bm{x}_{n}$, the observed variable $\bm{y}_{n}$ is drawn according to
@@ 3 GPLVMs and inference with GPLVMs
In order to expand our work on UQ to as wide a set of problems as possible, we explore the use of Random Fourier Features (RFF)–based GPs Rahimi and Recht, (2007). These GPs approximate an adopted kernel function by a finite-dimensional RFF mapping by
@@
where $\bm{\omega}_{j}$ are random frequencies, which according to Bochner’s theorem Rudin, (2017), are drawn from the power spectral density of the chosen kernel. The functions of interest shown in Fig. LABEL:Fig:GPLVM are then modeled by
@@
4 Uncertainty quantification: The theory
@@
5 Uncertainty quantification: The methods
@@
6 Uncertainty quantification: The insights from experiments
@@
During training, we drew $M=100$ samples from $p(\bm{X}|\bm{Y})$ using the variational distribution of $\bm{X}$. For each test sample, $\bm{y}_{o,*}$ and $\bm{X}^{(m)}$, we generated $L=100$ samples of $\bm{x}_{*}$ according to $\bm{x}_{*}^{(m,\ell)}\sim p(\bm{x}_{*}|\bm{y}_{o,*},\bm{X}^{(m)},\bm{Y})$. We used the generated training samples $\bm{X}^{(m)}$ to compute the epistemic and aleatoric uncertainties defined by (5) and (32), respectively. We conducted experiments where the number of drawn frequencies took values from the set $J\in\{10,25,50,100,200,500\}$. All the frequencies were obtained by sampling from the radial-basis-function kernel.
During evaluation, we computed the aleatoric and epistemic uncertainties as described in the previous section. In each trial, we assumed that one of the output variables, $y_{d,*}$, was missing and had to be predicted using the remaining variables. For example, to predict $y_{1,*}$, we used the observed values $y_{2,*}$, $y_{3,*}$, and $y_{4,*}$ to estimate the latent variable $\bm{x}_{*}$, which was then used to infer $y_{1,*}$. The boxplots in Figs. 1 and 2 summarize average aleatoric and epistemic variances computed on 200 test samples per trial, following training. Each boxplot is constructed from 50 such averages obtained across independent simulations. The circles denote outlier trials. The plots illustrate the consistency of uncertainty estimates and the variability across test instances.